the square root of 2k+1 =17
Answer by dolly(163)


Sq rt(2k+1) = 17
Squaring both sides,
2k+1 = 17^2
==> 2k+1 = 289
==> 2k = 269 - 1
==> 2k = 288
==> 2k/2 = 288/2
==> k = 144